a model rocket is launched from a raised platform at a speed of 144 ft./s it's height in feet is given by h(f)= -16t squared plus 144 t+20 (t=seconds after launch) After how many seconds does the object reach its maximum height
1
Expert's answer
2020-08-05T16:41:43-0400

"h(t)=-16t^2+144t+20"


"\frac{dh(t)}{dt}=0\to -32t+144=0\to t=144/32=4.5(s)"


After "4.5 (s)".